# Energy dynamics in a class of local random matrix Hamiltonians

### v2

@@ Abstract
Random matrix theory yields valuable insights into the universal features of quantum many-body chaotic systems. Although all-to-all interactions are traditionally studied, many interesting dynamical questions, such as transport of a conserved density, require a notion of spatially local interactions. We study the transport of the energy, the most basic conserved density, in few-body and 1D chains of nearest-neighbor random matrix terms that square to one. In the few-body but large local Hilbert space dimension case, we develop a mapping for the energy dynamics to a single-particle hopping picture. This allows for the computation of the energy density autocorrelators and an out-of-time-ordered correlator of the energy density. In the 1D chain, we numerically study the energy transport for a small local Hilbert space dimension. We also discuss the density of states throughout and touch upon the relation to free probability theory.
@@ I Introduction
Since the microscopic analytical treatment of specific strongly interacting systems (e.g. Hubbard model, QCD, quantum magnets, etc.) is essentially nonexistent, one can appeal to various phenomenological calculations to achieve analytical control. These include mean-field theory, large $N$ limits, conformal symmetry, or, for questions of “generic” or “statistical” properties, random matrix theory (RMT). Since Wigner and his successors’ work on the statistics of the energy levels of large nuclei [3], Hamiltonian random matrix theory has informed our understanding of chaos in the quantum mechanical setting [4]. This is particularly true in systems where no semi-classical limit is available [5, 6]. Traditionally, however, random matrix phenomenology treats a zero-dimensional system in the form of all-to-all interactions. This suffices to describe physics at the Heisenberg timescale (inverse of the typical many-body level-spacing) but for addressing certain dynamical questions such as the hydrodynamics of a local density raised in the previous paragraph, locality of interactions is necessary.
@@
It is interesting to ask if any of these tools can be extended to treat generic local energy conserving dynamics on the lattice. Then, one could try to prove that the presence of the most fundamental local conservation law, energy, leads to a universal emergent hydrodynamic description. Energy transport is also interesting because the globally conserved charge generates the dynamics itself. Unlike the case of spin transport, for energy transport the local charge density operators must fail to commute (otherwise the system would be non-interacting) making analytical treatment more challenging. One line of work treats the question of energy transport for the Sachdev-Ye-Kitaev (SYK) model [12, 13] and indeed diffusive energy transport has been established in 1D generalizations thereof [14, 15]. Here we are interested in such calculations for a random matrix generalization of a disordered locally interacting 1D quantum spin chain. For a pair of random matrices interacting in a strictly zero-dimensional manner, the energy dynamics were derived when the total density of states has a square-root dependence on energy at the edge of the spectrum (as in common in random matrix theories) [16]. Our work should be seen as complementary to that one, which will be referenced throughout.
@@ I.1 Local Haar-Ising random matrix model
where $h_{j}$ is a nearest-neighbor (NN) spin-spin like interaction term coupling sites $j,j+1$ (see Figure 1). We also refer to the interaction operators $h_{j}$ as “bonds”. Each local term is realized as $h_{j}=U_{j}\Lambda U^{\dagger}_{j}\otimes I_{\overline{j}}$ for which $\Lambda^{2}=1$, $\text{tr}(\Lambda)=0$, and $U_{j}$ are independent Haar random unitary matrices. The matrices $U_{j}$ and $\Lambda$ are $q^{2}\times q^{2}$, reflecting the NN locality of interaction, and $I_{\overline{j}}$ is an identity matrix acting on all sites besides $j,j+1$. We refer to $U_{j}\Lambda U^{\dagger}_{j}$ as Haar-Ising (HI) random matrices since their eigenvectors are Haar random and their eigenvalues are ±$\pm 1$ in equal number. We will consider both periodic (PBC) and open boundary conditions (OBC). A translation invariant version with $U_{j}=U$ could also be considered, though we focus on the disordered case in this paper. In the free probability literature, a self-adjoint random variable with this spectrum is referred to as a symmetric Bernoulli variable [23] and such matrices were also studied in [16] under the name “binary” random matrices.
The following features of the model Eq. (1) make it an interesting platform for studying generic out-of-equilibrium Hamiltonian dynamics. First, the local terms generalize Ising or spin 1/2 Heisenberg interactions of the form $\vec{S}_{j}\cdot\vec{S}_{j+1}$. However, we note that large $q$ does not correspond to large spin $S$. Secondly, the local terms have individually bounded eigenvalue spectra of ±$\pm 1$ even in a large $q$ limit, and so the Hamiltonian strictly obeys Lieb-Robinson bounds [24] independently of $q$ (for each arbitrarily large but finite $q$). Additionally, the maximum eigenvalue of the full Hamiltonian is bounded by $L$ independently of $q$, so there is no subtlety related to rescaling the energy with $q$ in order to make the system thermodynamically well-defined at large $q$. Thirdly, unlike a chain of local terms drawn from the Gaussian Unitary Ensemble (GUE), the terms we study are not individually quantum chaotic since they have massively degenerate eigenvalue spectra. This is interesting from a many-body quantum chaos point of view, since the putative chaoticity of the model would only arise through the interplay of the terms. One can imagine that the dynamics of the spectral form factor (SFF) [25] would differ from a theory with locally chaotic terms, where (at least in discrete time) there is a transition from the early time behavior $t^{L}$ characteristic of $L$ non-interacting chaotic systems, to the universal late-time ramp $t$ [26]. Considering two interacting terms in particular, perhaps in the GUE case there is an early time $t^{2}$ behavior of the connected SFF, whereas for the HI case there would be no such regime. We reserve a study of chaos for future work.
@@ I.2 Energy dynamics
Anomalous energy transport is also possible in non-integrable systems; for example, in the presence of kinetic constraints such as PXP dynamics, Kardar-Parisi-Zhang superdiffusion was found [29]. If the system is disordered, one might anticipate localization phenomena, and, for example, a sufficiently disordered XYZ chain exhibits sub-diffusion [30] of energy at high temperature. However, despite our $q=2$ Haar-Ising random matrix spin chain being in some sense maximally disordered, we do not find any evidence of energy localization, and instead find numerical evidence of a slow convergence to diffusion. This is consistent with the level spacing statistics being that of a GUE and not Poisson (which would correspond to many-body localization [31]).
@@ I.3 Summary and organization of results
In Sec. II we study two Haar-Ising terms overlapping a single site on a three site lattice. We show that infinite-temperature equal-time correlators of the local terms are either one or zero in the large $q$ limit, which also proves that the two terms are asymptotically free in the language of free probability [43], despite their eigenvectors not being related by a full Haar random unitary. With this, we establish a mapping to a free particle hopping on a 1D lattice and use it to analytically compute the density of states, energy-energy correlators, and an OTOC at infinite temperature. We also calculate the energy-energy correlator at finite temperature.
In Sec. III we generalize this mapping to treat $z$ overlapping Haar-Ising terms in a situation where none commute. Subject to a certain conjecture, we map the dynamics to a particle hopping on the Bethe lattice with coordination number $z$ and then compute the density of states and energy-energy correlators.
Ultimately, we would like to treat the energy transport analytically in the fully extended 1D HI chain. Since the introduction of true locality (in the sense of some terms commuting and some not) makes the analysis more difficult, we initiate the study by numerically calculating the energy transport for a $q=2$ chain of up to $L=22$ spins. We scrutinize the emergence of energy diffusion in the system by comparing against the solution to the diffusion equation on the same finite lattice, and find that while the autocorrelator quickly converges to that form, the spatially distant two-point functions are slower to converge.
In Sec. V we also address the simpler question of the density of states of various few-body and extended HI chains and make the relation of our previous analysis to free probability theory more explicit. We numerically identify some situations where free probability can be naively applied to compute the density of states, and some where it cannot.
We conclude in Sec. VI with some remarks about how we may generalize the single-particle mapping to treat the extended HI RMT chain as well as some remarks about chaos.
@@ II.1 Alternating correlators vanish
Given such a word of the local terms, we can apply the algebraic relations $h_{i}^{2}=1$ to fully reduce the word. If it cannot be reduced to the identity operator, then we call it irreducible. The string of indices of (fully reduced) irreducible words must be of one of the following forms: $(01)^{m}$$(10)^{m}$$1(01)^{m}$$0(10)^{m}$. Since we are interested in traces of the random matrix terms, the first two strings are equivalent, while the latter two can be further reduced to $1$ and 00, respectively, due to cyclicity of the trace. Those latter two vanish since $\braket{h_{i}}=0$ by construction. Regarding the first type, we assert the non-trivial claim that the alternating words $(01)^{m}$ also vanish in the large $q$ limit, i.e.
@@
This is proven by invoking the Weingarten calculus for averaging over the unitary group.
@@
Now, an important fact is that the odd free cumulants of a symmetric Bernoulli variable and therefore the HI random matrices are zero [23]. Therefore, if $\mu$ or $\nu$ contain at least one cycle of odd length, then that term will not contribute to the sum Eq. (15). We will show that consequently, there are no non-zero $O(q^{0})$ terms (nor any $O(q^{-1})$ terms) in that sum. First, if $m$ is odd, both $\mu$ and $\nu$ necessarily contain an odd length cycle so the correlator is identically zero. If $m$ is even, however, the following argument shows that $g(\mu,\nu)\geq 1$. Because $\mu$ and $\nu$ contain only even length cycles, we have
@@
with no prefactor. We show in Fig. 3(a) that this scaling with $q$ can be seen numerically already for $q=8,10,12,\dots$ in this $m=2$ case  [^1] For larger $m$, e.g. $m=4$, the realization-to-realization fluctuations appear to be significantly larger (e.g. 80×80$\times$80 × larger) than the average and so the scaling of the true average is difficult to see numerically..
@@ II.2 Density of states
We can map the counting of reducible strings to counting trajectories of a particle on an infinite 1D lattice. A general bitstring (reducible or not) maps to a trajectory as follows. Each bit in the string (starting from the right without loss of generality) corresponds to the particle hopping once. Whether the hop is to the right or to the left, however, depends upon the parity of the current site. If the particle sits at an even site (blue lattice points in Fig. 2(b)), a 00 moves it to the right and a $1$ moves it left, but the opposite holds if the particle is at an odd site (orange lattice points in the same figure). This is set up so that reducible strings correspond to walks that return to their starting location, e.g. the string $00$ corresponds to hopping once right and then left. Note that the lattice points are in one-to-one correspondence with fully irreducible words.
@@
where |$\ket{x}$ is a single particle sitting at site $x$. Since |$\ket{0}$ is the particle sitting at the origin, we have for example
@@
One can then see that the number of reducible words/walks of length $n$ is given by the amplitude to return. Therefore we have the correspondence $\braket{H^{n}}=\braket{0}{\Delta^{n}}{0}$. These vanish for odd $n$, and, for $n=2k$, the sequence is given by the central binomial coefficient
@@ II.3 Local energy correlator
Recalling the mapping to the 1D lattice hopping, the string $a0b0$ corresponds to the following process when reading from right to left. Starting at the origin, we first hop right, then undergo an arbitrary walk of length $m$. The particle has undergone the walk dictated by the string $b0$, and it will now be sitting at some lattice point $-m+1\leq x\leq m+1$ along the line. The next step must be 00. Depending upon the parity of $x$ (equivalently whether the corresponding irreducible string ends in a 00 or a $1$), it will hop left or right. Finally, it will undergo another walk of length $n$ dictated by $a$ which starts at $x\pm 1$ and ends at the origin $x=0$.
@@
$V=\sum_{x\ \text{odd}}\ket{x-1}\bra{x}+\sum_{x\ \text{even}}\ket{x+1}\bra{x}.$ | . (29)
This hops a particle right or left depending on the parity of the site. The strings of reducible words $a0b0$ are then counted by the amplitude $\braket{0}{\Delta^{n}V\Delta^{m}V}{0}$, i.e. we have established the correspondence in the large $q$ limit,
$\braket{H^{n}hH^{m}h}=\braket{0}{\Delta^{n}V\Delta^{m}V}{0}$ (30)
@@
|$\ket{k}=\frac{1}{\sqrt{N}}\sum_{x}e^{ikx}\ket{x},\quad k\in\frac{2\pi n}{N}$ (32)
where the set $n\in\{0,1,2,\dots,N-1\}$.
One more useful fact is the following translation invariance  [^2] This is not exactly trivial; but follows by inserting $I=T_{x}^{\dagger}T_{x}$ where $T_{x}$ translates by $x$ sites: $\braket{0}{V(t)V(0)}{0}=\braket{0}{T^{\dagger}_{x}T_{x}V(t)T_{x}^{\dagger}T_{x}V(0)T_{x}^{\dagger}T_{x}}{0}.$ By using that $\Delta$ is translation invariant, we have $\braket{0}{V(t)V(0)}{0}=\braket{x}{(T_{x}VT^{\dagger}_{x})(t)(T_{x}VT^{\dagger}_{x})(0)}{x}.$ If $x$ is even, then $T_{x}VT^{\dagger}_{x}=V$. If $x$ is odd, then $T_{x}VT^{\dagger}_{x}=V^{\prime}$ where $V^{\prime}$ now hops an even site particle to the left which is the opposite action of $V$. By symmetry of the original RMT model under $h_{0}\leftrightarrow h_{1}$, we still have the claimed equality for odd $x$.
$\braket{0}{V(t)V(0)}{0}=\braket{x}{V(t)V(0)}{x}\quad\forall x.$ (33)
@@
$\operatorname{tr}_{1}[V(t)V(0)]=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{x}\braket{x}{V(t)V(0)}{x},$ (34)
@@
$C(t)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{k}\braket{k}{V(t)V(0)}{k},$ (35)
@@
[…]
@@
$\braket{k}{e^{-i\Delta t}}{p}=e^{-i\epsilon_{k}t}\delta_{k,p},$ (37)
@@
$\braket{k}{V}{p}=\cos(k)\delta_{k,p}+i\sin(k)\delta_{k,p+\pi}.$ (38)

### v3

@@ Abstract
Random matrix theory yields valuable insights into the universal features of quantum many-body chaotic systems. Although all-to-all interactions are traditionally studied, many interesting dynamical questions, such as transport of a conserved density, require a notion of spatially local interactions. We study the transport of the energy, the most basic conserved density, in few-body and 1D chains of nearest-neighbor random matrix terms that square to one, which we dub the “Haar-Ising” local random matrix model. In the few-body but large local Hilbert space dimension case, we develop a mapping for the energy dynamics to a single-particle hopping picture. This allows for the computation of the energy density autocorrelators and an out-of-time-ordered correlator of the energy density. In the 1D chain, we numerically study the energy transport for a small local Hilbert space dimension. Throughout, we discuss the density of states and touch upon the relation to free probability theory.
@@ I Introduction
Since the microscopic analytical treatment of specific strongly interacting systems (e.g. Hubbard model, QCD, quantum magnets, etc.) is essentially nonexistent, one can appeal to various phenomenological calculations to achieve analytical control. These include mean-field theory, large $N$ limits, conformal symmetry, or, for questions of “generic” or “statistical” properties, random matrix theory (RMT). Since Wigner and his successors’ work on the statistics of the energy levels of large nuclei [3], Hamiltonian random matrix theory has informed our understanding of chaos in the quantum mechanical setting [4]. This is particularly true in systems where no semi-classical limit is available [5, 6].
Standard random matrix theory phenomenology treats a zero-dimensional system in the form of all-to-all interactions. Various aspects of locality have, however, been incorporated in the literature. These include various types of banded random matrices [7] whose matrix elements decay away from the main diagonal, Wegner’s $n$-orbital model describing disordered hopping in a large $n$ limit [8], or Rosenzweig-Porter models whose diagonal elements are stronger than off-diagonal, but which still maintain all-to-all interaction [9]. Local interactions also give rise to sparse Hamiltonians, which have been treated [10], along with Sachdev-Ye-Kitaev (SYK) models which can also be viewed as sparse random matrices and are discussed later in this section. Certain manifestations of interactions (through correlations) have also been introduced in the RMT literature [11].
Many of these models succeed in describing universal physics at the Heisenberg timescale (inverse of the typical many-body level-spacing) of various interacting chaotic systems or disordered non-interacting ones (e.g., level statistics in the localized and delocalized phases and at the critical point of the Anderson localization transition). Our primary motivations are somewhat different. We are interested primarily in dynamics of a generic Hamiltonian system subject to geometric locality of interaction, for which one can treat questions of hydrodynamics of a local conserved density. Such questions also involve “early-time” dynamics, e.g. $t\sim O(L^{\alpha})$ in a system of $L$ local subsystems, as opposed to the Heisenberg time-scale $t\sim e^{O(L)}$ that is relevant for level statistics.
@@
It is interesting to ask if any of these tools can be extended to treat generic local energy conserving dynamics on the lattice. Then, one could try to prove that the presence of the most fundamental local conservation law, energy, leads to a universal emergent hydrodynamic description. Energy transport is also interesting because the globally conserved charge generates the dynamics itself. Unlike the case of spin transport, for energy transport the local charge density operators must fail to commute (otherwise the system would be non-interacting) making analytical treatment more challenging. One line of work treats the question of energy transport for the Sachdev-Ye-Kitaev (SYK) model [17, 18] and indeed diffusive energy transport has been established in 1D generalizations thereof [19, 20]. Here we are interested in such calculations for a random matrix generalization of a disordered locally interacting 1D quantum spin chain. For a pair of random matrices interacting in a strictly zero-dimensional manner, the energy dynamics were derived when the total density of states has a square-root dependence on energy at the edge of the spectrum (as is common in random matrix theories) [21]. Our work should be seen as complementary to that one, which will be referenced throughout.
@@ I.1 Local Haar-Ising random matrix model
where $h_{j}$ is a nearest-neighbor (NN) spin-spin like interaction term coupling sites $j,j+1$ (see Figure 1). We also refer to the interaction operators $h_{j}$ as “bonds”. Each local term is realized as $h_{j}=U_{j}\Lambda U^{\dagger}_{j}\otimes I_{\overline{j}}$ for which $\Lambda^{2}=1$, $\text{tr}(\Lambda)=0$, and $U_{j}$ are independent Haar random unitary matrices. The matrices $U_{j}$ and $\Lambda$ are $q^{2}\times q^{2}$, reflecting the NN locality of interaction, and $I_{\overline{j}}$ is an identity matrix acting on all sites besides $j,j+1$. We refer to $U_{j}\Lambda U^{\dagger}_{j}$ as Haar-Ising (HI) random matrices since their eigenvectors are Haar random and their eigenvalues are $\pm 1$ in equal number. We will consider both periodic (PBC) and open boundary conditions (OBC). A translation invariant version with $U_{j}=U$ could also be considered, though we focus on the disordered case in this paper. In the free probability literature, a self-adjoint random variable with this spectrum is referred to as a symmetric Bernoulli variable [28] and such matrices were also studied in [21] under the name “binary” random matrices.
The following features of the model Eq. (1) make it an interesting platform for studying generic out-of-equilibrium Hamiltonian dynamics. First, the local terms generalize Ising or spin 1/2 Heisenberg interactions of the form $\vec{S}_{j}\cdot\vec{S}_{j+1}$. However, we note that large $q$ does not correspond to large spin $S$. Secondly, the local terms have individually bounded eigenvalue spectra of $\pm 1$ even in a large $q$ limit, and so the Hamiltonian strictly obeys Lieb-Robinson bounds [29] independently of $q$ (for each arbitrarily large but finite $q$). Additionally, the maximum eigenvalue of the full Hamiltonian is bounded by $L$ independently of $q$, so there is no subtlety related to rescaling the energy with $q$ in order to make the system thermodynamically well-defined at large $q$.
Thirdly, unlike a chain of local terms drawn from the Wigner-Dyson ensembles (GUE, GOE or GSE), the terms we study are not individually quantum chaotic since they have massively degenerate eigenvalue spectra. This is interesting from a many-body quantum chaos point of view, since the putative chaoticity of the model would only arise through the interplay of the terms. One can imagine that the dynamics of the spectral form factor (SFF) [30] would differ from a theory with locally chaotic terms, where (at least in discrete time) there is a transition from the early time behavior $t^{L}$ characteristic of $L$ non-interacting chaotic systems, to the universal late-time ramp $t$ [31]. Considering two interacting terms in particular, perhaps in the GUE case there is an early time $t^{2}$ behavior of the connected SFF, whereas for the HI case there would be no such regime. We reserve a study of chaos for future work. We also note that quantum chaos through the perspective of the eigenstate thermalization hypothesis (ETH) was studied in the situation of local RMT Hamiltonians [32, 33].
@@ I.2 Energy dynamics
Anomalous energy transport is also possible in non-integrable systems; for example, in the presence of kinetic constraints such as PXP dynamics, Kardar-Parisi-Zhang superdiffusion was found [36]. If the system is disordered, one might anticipate localization phenomena, and, for example, a sufficiently disordered XYZ chain exhibits sub-diffusion [37] of energy at high temperature. However, despite our $q=2$ Haar-Ising random matrix spin chain being in some sense maximally disordered, we do not find any evidence of energy localization, and instead find certain numerical signatures of diffusion. The lack of localization is consistent with the level spacing statistics being that of a GUE and not Poisson (which would correspond to many-body localization [38]).
@@ I.3 Summary and organization of results
In Sec. II we study two Haar-Ising terms overlapping a single site on a three site lattice. We show that infinite-temperature equal-time correlators of the local terms are either one or zero in the large $q$ limit, which also proves that the two terms are asymptotically free in the language of free probability [50], despite their eigenvectors not being related by a full Haar random unitary. The proof also follows from a recent result in free probability [26]. With this, we establish a mapping to a free particle hopping on a 1D lattice and use it to analytically compute the density of states, energy-energy correlators, and an OTOC at infinite temperature. We also calculate the energy-energy correlator at finite temperature.
In Sec. III we generalize this mapping to treat $z$ overlapping Haar-Ising terms in a situation where none commute. We map the dynamics to a particle hopping on the Bethe lattice with coordination number $z$ and then compute the density of states and energy-energy correlators. We will see that in all of these cases $z\geq 2$, the dynamics of the local energy density does not follow a diffusive form. This is due to the few-body nature of the Hamiltonian, which is local in the sense that it is more sparse than a full random matrix, but not geometrically local as in the full extended chain.
Ultimately, we would like to treat the energy transport analytically in this full extended 1D HI chain. Since the introduction of true locality (in the sense of some terms commuting and some not) makes the analysis more difficult, we initiate the study by numerically calculating the energy transport for a $q=2$ chain of up to $L=22$ spins. We scrutinize the emergence of energy diffusion in the system by comparing against the solution to the diffusion equation on the same finite lattice, and find that while the autocorrelator quickly converges to that form, the spatially distant two-point functions are slower to converge.
In Sec. V we also address the simpler question of the density of states of various few-body and extended HI chains and make the relation of our previous analysis to free probability theory more explicit. We numerically identify some situations where standard techniques from free probability can be naively applied to compute the density of states, and some where they cannot.
We conclude in Sec. VI with some comments about how we may generalize the single-particle mapping, including numerical evidence that only the reducible words survive a large $q$ limit more generally, as well as some remarks about chaos.
@@ II.1 Alternating correlators vanish
Given such a word of the local terms, we can apply the algebraic relations $h_{i}^{2}=1$ to fully reduce the word. If it cannot be reduced to the identity operator, then we call it irreducible. The string of indices of (fully reduced) irreducible words must be of one of the following forms: $(01)^{m}$$(10)^{m}$$1(01)^{m}$$0(10)^{m}$. Since we are interested in traces of the random matrix terms, the first two strings are equivalent, while the latter two can be further reduced to 0 and $1$, respectively, due to cyclicity of the trace. Those latter two vanish since $\braket{h_{i}}=0$ by construction. Regarding the first type, we assert the non-trivial claim that the alternating words $(01)^{m}$ also vanish in the large $q$ limit, i.e.
@@
This is proven by invoking the Weingarten calculus for averaging over the unitary group, which we find instructive to work out in the remainder of this section. The result also, however, follows from recent work in free probability theory [26, 51].
@@
Now, an important fact is that the odd free cumulants of a symmetric Bernoulli variable and therefore the those of the HI random matrices are zero [28]. Therefore, if $\mu$ or $\nu$ contain at least one cycle of odd length, then that term will not contribute to the sum Eq. (15). We will show that consequently, there are no non-zero $O(q^{0})$ terms (nor any $O(q^{-1})$ terms) in that sum. First, if $m$ is odd, both $\mu$ and $\nu$ necessarily contain an odd length cycle so the correlator is identically zero. If $m$ is even, however, the following argument shows that $g(\mu,\nu)\geq 1$. Because $\mu$ and $\nu$ contain only even length cycles, we have
@@
with no prefactor. We show in Fig. 3(a) that this scaling with $q$ can be seen numerically already for $q=8,10,12,\dots$ in this $m=2$ case  [^1] For larger $m$, e.g. $m=4$, the realization-to-realization fluctuations appear to be significantly larger (e.g. $80\times$ larger) than the average and so the scaling of the true average is difficult to see numerically..
@@ II.2 Density of states
We can map the counting of reducible strings to counting trajectories of a particle on an infinite 1D lattice. A general bitstring (reducible or not) maps to a trajectory as follows. Each bit in the string (starting from the right without loss of generality) corresponds to the particle hopping once. Whether the hop is to the right or to the left, however, depends upon the parity of the current site. If the particle sits at an even site (blue lattice points in Fig. 2(b)), a 0 moves it to the right and a $1$ moves it left, but the opposite holds if the particle is at an odd site (orange lattice points in the same figure). This is set up so that reducible strings correspond to walks that return to their starting location, e.g. the string $00$ corresponds to hopping once right and then left. Note that the lattice points are in one-to-one correspondence with fully irreducible words.
@@
[…]
One can then see that the number of reducible words/walks of length $n$ is given by the amplitude to return. Therefore we have the correspondence $\braket{H^{n}}=\braket{0|\Delta^{n}|0}$. These vanish for odd $n$, and, for $n=2k$, the sequence is given by the central binomial coefficient
@@ II.3 Local energy correlator
Recalling the mapping to the 1D lattice hopping, the string $a0b0$ corresponds to the following process when reading from right to left. Starting at the origin, we first hop right, then undergo an arbitrary walk of length $m$. The particle has undergone the walk dictated by the string $b0$, and it will now be sitting at some lattice point $m+1\leq x\leq m+1$ along the line. The next step must be 0. Depending upon the parity of $x$ (equivalently whether the corresponding irreducible string ends in a 0 or a $1$), it will hop left or right. Finally, it will undergo another walk of length $n$ dictated by $a$ which starts at $x\pm 1$ and ends at the origin $x=0$.
@@
[…]
This hops a particle right or left depending on the parity of the site. The strings of reducible words $a0b0$ are then counted by the amplitude $\braket{0|\Delta^{n}V\Delta^{m}V|0}$, i.e. we have established the correspondence in the large $q$ limit,
$\braket{H^{n}hH^{m}h}=\braket{0|\Delta^{n}V\Delta^{m}V|0}$ (30)
@@
[…]
where $n\in\{0,1,2,\dots,N-1\}$.
One more useful fact is the following translation invariance  [^2] This is not exactly trivial; but follows by inserting $I=T_{x}^{\dagger}T_{x}$ where $T_{x}$ translates by $x$ sites: $\braket{0|V(t)V(0)|0}=\braket{0|T^{\dagger}_{x}T_{x}V(t)T_{x}^{\dagger}T_{x}V(0)T_{x}^{\dagger}T_{x}|0}.$ By using that $\Delta$ is translation invariant, we have $\braket{0|V(t)V(0)|0}=\braket{x|(T_{x}VT^{\dagger}_{x})(t)(T_{x}VT^{\dagger}_{x})(0)|x}.$ If $x$ is even, then $T_{x}VT^{\dagger}_{x}=V$. If $x$ is odd, then $T_{x}VT^{\dagger}_{x}=V^{\prime}$ where $V^{\prime}$ now hops an even site particle to the left which is the opposite action of $V$. By symmetry of the original RMT model under $h_{0}\leftrightarrow h_{1}$, we still have the claimed equality for odd $x$.
$\braket{0|V(t)V(0)|0}=\braket{x|V(t)V(0)|x}\quad\forall x.$ (33)
@@
$\operatorname{tr}_{1}[V(t)V(0)]=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{x}\braket{x|V(t)V(0)|x},$ (34)
@@
$C(t)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{k}\braket{k|V(t)V(0)|k},$ (35)
@@
$C(t)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{kp}\braket{k|e^{-i\Delta t}|k}|\braket{k|V|p}|^{2}\braket{p|e^{i\Delta t}|p}.$ (36)
@@
$\braket{k|e^{-i\Delta t}|p}=e^{-i\epsilon_{k}t}\delta_{k,p},$ (37)
@@
$\braket{k|V|p}=\cos(k)\delta_{k,p}+i\sin(k)\delta_{k,p+\pi}.$ (38)